OpenStudy (anonymous):

complete the square ? x2 + 22x = 0

OpenStudy (amistre64):

tell me what you think it is and Ill tell you what I think it is ;)

OpenStudy (anonymous):

im sorry , but can put an example of how im suppose to set it up .

OpenStudy (amistre64):

sure: we see that 22 sitting all nice and pretty there. we want to use it. 22. now lets split it in half,,,right in two. 22/2 = 11 we got our number that we want to play with. so lets multiply it by itself: 11 * 11 = 121 and thats the answer we want. 121

OpenStudy (anonymous):

22/2=11 11^2=121 x^2+11x+121 (x+11)^2